# FADPNet: Frequency-Aware Dual-Path Network for Face Super-Resolution

### v1

@@ Abstract
Face super-resolution (FSR) under limited computational costs remains an open problem. Existing approaches typically treat all facial pixels equally, resulting in suboptimal allocation of computational resources and degraded FSR performance. CNN is relatively sensitive to high-frequency facial features, such as component contours and facial outlines. Meanwhile, Mamba excels at capturing low-frequency features like facial color and fine-grained texture, and does so with lower complexity than Transformers. Motivated by these observations, we propose FADPNet, a Frequency-Aware Dual-Path Network that decomposes facial features into low- and high-frequency components and processes them via dedicated branches. For low-frequency regions, we introduce a Mamba-based Low-Frequency Enhancement Block (LFEB), which combines state-space attention with squeeze-and-excitation operations to extract low-frequency global interactions and emphasize informative channels. For high-frequency regions, we design a CNN-based Deep Position-Aware Attention (DPA) module to enhance spatially-dependent structural details, complemented by a lightweight High-Frequency Refinement (HFR) module that further refines frequency-specific representations. Through the above designs, our method achieves an excellent balance between FSR quality and model efficiency, outperforming existing approaches.
@@ I Introduction
Face super-resolution (FSR), also known as face hallucination, focuses on reconstructing high-resolution (HR) facial images from low-resolution (LR) inputs. FSR plays a critical role in real-world scenarios where LR faces may appear in surveillance videos, historical archives, or mobile imaging. Enhancing facial image resolution is essential for improving downstream tasks such as face recognition [1], face alignment [2], and facial attribute analysis [3].
Figure 1: (a) Spectral energy distribution of features extracted by Mamba and CNN architectures, where Mamba exhibits concentrated low-frequency responses compared to CNN’s high-frequency sensitivity. Color intensity corresponds to normalized energy values. (b) Efficiency trade-offs on CelebA test set [4], demonstrating our method’s optimal balance between PSNR, Params, and inference speed.
@@
One fundamental challenge in FSR lies in the long-tailed distribution of image frequency components [5]. High-frequency parts, such as the eyes, nose, mouth, and facial contours, account for a small portion of the facial area but require disproportionate modeling capacity due to their high variability. These regions are particularly sensitive to identity, illumination, and expression changes, making them significantly harder [6] to reconstruct than low-frequency areas such as skin, which contribute little to identity-specific features.. This imbalance results in a performance bottleneck, where failure to recover high-frequency content leads to perceptual degradation. However, most existing methods treat all pixels equally, ignoring the skewed distribution of visual complexity and importance. This uniform processing strategy leads to inefficient resource usage and suboptimal reconstruction quality. While some methods attempt to separate frequency components [7, 8], they often struggle to maintain a balanced representation across global and local facial structures.
To address these limitations, we aim to adaptively distinguish and process high- and low-frequency information. Meanwhile, we wonder about a dual-path feature fusion to help resolve the imbalance between local and global modeling that arises in previous frequency-aware methods, ultimately enabling more efficient and accurate FSR. As shown in Fig. 1(a), we present the spectral energy maps generated by CNN [9], and Mamba [10]. Interestingly, Mamba exhibits broader responses in the low-frequency regions of facial images, whereas CNN tends to be more sensitive to high-frequency components. Moreover, compared to the Transformer [11] framework, Mamba achieves global modeling through scanning mechanisms, thus keeping the computational complexity within a linear range and avoiding the expensive quadratic cost. These observations lead us to consider an alternative design: could we develop a two-branch architecture that allows Mamba and CNN to each specialize in the frequency domains they are most suited for, while integrating their outputs to better balance global and local facial structure representation?
@@
Based on the above observations and analysis, we propose a Frequency-Aware Dual-Path Network (FADPNet), which allocates distinct submodules to specialize in different frequency domains. In particular, we introduce the Low-Frequency Enhancement Block (LFEB) and High-Frequency Enhancement Block (HFEB) to extract and refine, respectively, stable identity-preserving low-frequency features and fine-grained, structure-critical high-frequency details. The LFEB integrates an Attentive State Space Block (ASSB) and a Squeeze-and-Excitation Block (SEB) to enhance low-frequency features. At the same time, the HFEB combines a High-Frequency Refinement (HFR) module and a Depthwise Position-aware Attention (DPA) module to refine high-frequency details. These two modules integrate complementary modeling strategies—global and local—to better capture the intrinsic characteristics of facial components across different frequency levels. By decomposing and processing features based on their frequency significance, our approach not only enhances facial fidelity but also improves computational efficiency by focusing resources on perceptually important regions. As shown in Fig. 1(b), our proposed FADPNet achieves superior efficiency and reconstruction quality compared to existing FSR methods. In summary, the main contributions are as follows:
@@ II-A Face Super-Resolution
Face super-resolution [12] has made notable strides with the advancement of deep learning. Early approaches [13, 8, 14, 15] leverage explicit facial priors to guide reconstruction. For instance, FSRNet [13] incorporated facial landmarks and parsing maps to enforce structural consistency, while FAN [8] applied a face alignment network with progressive training to enhance realism. DIC [14] introduced an iterative refinement loop combining facial components and landmark predictions. To improve the representation of facial spatial features, Hu et al. [15] first utilized 3D shape priors in the network to better preserve sharp facial spatial structures.
However, prior-based methods depend heavily on accurate prior estimation, which is unreliable at extremely LR scenes and adds computational overhead. To mitigate this, recent works have adopted attention-based [16] or data-driven strategies [17]. SPARNet [16] used spatial attention to focus on key facial regions. SISN [18] decoupled attention into structural and textural streams. AD-GNN [19] combined spatial attention with graph-based message passing to model feature dependencies. LAAT [20] introduced local-aware refinement within a Transformer framework to enhance detail, while SCTANet [21] jointly modeled spatial and channel attention. SFMNet [22] employed parallel frequency-spatial branches for multi-scale feature learning. WFEN [23] used wavelet-based encoding to mitigate downsampling artifacts. However, existing FSR methods apply uniform processing across spatial regions, overlooking the varying complexity of frequency components. In contrast, our method introduces a frequency-aware dual-path design that adaptively handles high- and low-frequency features at multiple scales, leading to more robust and detail-preserving reconstruction under LR conditions.
@@ II-B Long-Range Modeling
Modeling long-range dependencies is vital for FSR, as facial structures span distant spatial regions. Transformer-based methods [20, 22, 24] have demonstrated strong global modeling via self-attention. For example, LAAT [20] enhanced fine-grained facial modeling via self-refinement in a Transformer framework, improving detail restoration. SFMNet [22] adopted spatial and frequency branches to jointly capture global and local features across diverse receptive fields. CTCNet [25] further enhanced FSR by fusing multi-scale global representations extracted from Transformers.
However, vision Transformers [26, 25, 27] suffer from quadratic complexity, limiting their scalability. This challenge has prompted exploration of alternative long-range modeling with lower computational cost. Structured State-Space Models (SSMs)[28] have gained attention for their linear-time global modeling capability. Notably, Mamba [10] introduced a selective scan mechanism for efficient long-sequence processing while preserving strong global context, and has been successfully adapted to visual tasks [29, 30]. Building on this, MambaIR [31] extended Mamba to image restoration, alleviating issues such as local pixel forgetting and redundant channel representations. Variants like MambaLLIE [32] further demonstrate its adaptability to various low-level vision tasks.
Despite its efficiency, Mamba inherits a causal bias from its sequence-oriented design, which conflicts with the non-causal nature of super-resolution [31] and limits spatial modeling. To address this, MambaIRv2 [33] introduced direction-agnostic scanning for richer context modeling with linear scalability. Unlike existing Mamba-based methods [31, 32, 33], our approach incorporates a customized non-causal state-space backbone into the low-frequency branch, explicitly designed for face super-resolution to better preserve structural consistency and global facial context.
@@ II-C Frequency-Aware Super-Resolution
Frequency-aware SR methods have emerged as an effective paradigm for enhancing structure, particularly in edges and textures [34, 35, 36]. By decomposing image content into distinct frequency bands, these methods enable neural networks to more precisely model the unique characteristics of each frequency component [37]. This design aligns with the well-established observation that low-frequency components mainly represent smooth and homogeneous regions, whereas high-frequency components capture fine-grained details and edge features essential for perceptual quality [38].
Explicit transformation methods decompose the image reconstruction process into frequency-specific operations. For instance, wavelet-based methods like SRCliqueNet [38] applied a discrete wavelet transform to separate LR images into sub-bands, which are then processed by specialized subnetworks to sharpen high-frequency components while preserving low-frequency structures. Similarly, FreqNet [35] operated in the discrete cosine transform domain, predicting HR frequency coefficients through learnable spectral decoders. In contrast, implicit methods achieve frequency awareness through architectural design. SRDN [39] used high-pass constrained convolutions and residual connections to emphasize high-frequency propagation, while FSN [34] adopted Octave Convolution to split features into frequency-aware streams. AFD [40] estimated feature drifting in the DCT frequency domain, adaptively enhances features. However, the above methods either concentrate solely on high-frequency enhancement [41] or fail to model the interaction between global low-frequency structures and local high-frequency details. To address this, we propose a dual-branch framework that separately models frequency-specific features, improving both reconstruction quality and computational efficiency.
@@ III-A Overview of FADPNet
The proposed network architecture, as illustrated in Fig. 2, implements a hierarchical frequency-aware processing framework. It commences with an input image $\mathbf{I}_{LR}\in\mathbb{R}^{3\times H\times W}$ from which shallow facial features ${\mathbf{f}_{1}}\in\mathbb{R}^{C\times H\times W}$ are extracted. Our architecture adopts a three-level U-Net [42] structure, where each resolution stage includes a cascaded basic block that decomposes features into high- and low-frequency components for separate processing. The High-Frequency Enhancement Block (HFEB) restores fine-grained details using convolutions, Residual Blocks (RB), a High-Frequency Refinement (HFR) module, and a Depthwise Position-Aware Attention (DPA) module. The Low-Frequency Enhancement Block (LFEB) employs a dual-branch design: the Squeeze-and-Excitation Block (SEB) enhances local identity cues, while the Attentive State Space Block (ASSB) captures global facial structure. Outputs are fused and refined through a Feed-Forward Network (FFN). To reduce spatial misalignment caused by downsampling and upsampling, we introduce an offset-based warping module that aligns coarse structures with fine-level details, improving cross-scale consistency. Finally, features from all levels are fused and dimensionality reduction to reconstruct an HR face image $\mathbf{I}_{SR}\in\mathbb{R}^{3\times H\times W}$.
@@ III-B Low-Frequency Enhancement Block (LFEB)
Our Low-Frequency Enhancement Block (LFEB) is designed to model low-frequency facial features, such as smooth skin regions on the forehead, cheeks, and chin, which exhibit gradual intensity variations and contain minimal identity-specific details. Specifically, given a low-frequency feature $\mathbf{F}_{L}\in\mathbb{R}^{C\times H\times W}$ extracted from the frequency separation module. First, LFEB begins with a layer normalization for stable training, followed by a dual-branch structure: the Squeeze-and-Excitation Block (SEB) [43] for adaptively calibrating feature channels to improve the focus on important features, while the Attentive State Space Block (ASSB) captures global facial structure. A learnable fusion scheme adaptively balances two feature streams, guided by scaling parameters optimized during training. The fused output is further processed by a Feed-Forward Network (FFN) [10] and combined with the block’s input via a residual connection. This design enhances identity-salient, low-frequency attributes such as coarse facial shape and alignment while preserving photorealistic appearance.
Figure 3: The architecture of Attentive State Space Block (ASSB)
Figure 4: The architecture of (a) High-Frequency Refinement (HFR) (b) Depthwise Position-aware Attention (DPA)
@@ III-B1 Attentive State Space Block (ASSB)
where ${r}$ is the inner rank, ${T}$ is the number of prompts and ${d}$ is the number of hidden states in Mamba. Instance-specific prompt matrix is governed by a routing mechanism employing $\operatorname{gumbel-softmax}$ [44], which generates ${L}$ instance-specific prompts ${\mathbf{P}_{\text{m}}\in\mathbb{R}^{L\times T}}$ dynamically associated input pixels with relevant prompts from the pool:
@@
where ${\mathrm{W}_{\text{p}}}$ denotes the linear layer, $\operatorname{LogSoftmax}$ refers to the logarithm of the softmax function, and gumbel-softmax is a gumbel softmax [44] operator. Then, the Semantic Guided Neighboring (SGN) unfolds this 2D feature $\mathbf{F}_{2}$ into 1D sequences for SSE. SGN-Unfold mechanism redefines the traditional spatial-to-sequential transformation by dynamically reorganizing pixels into 1D sequences based on semantic similarity, effectively mitigating the long-range decay inherent to existing models [33]. Notably, the semantic information that the mechanism relies on during this process stems from the prompt matrix $\mathbf{P}_{\text{m}}$:
@@
${{y_{i}=(\mathbf{C}+\mathbf{P})h_{i}+\mathbf{D}x_{i}}}$ (8)
Finally, an SGN-fold is applied as the inverse operation, reconstructing the semantically sorted sequence back into the original spatial layout to form the 2D feature. This reversion is guided by the previously computed reverse index, ensuring accurate spatial alignment. The restored feature is then passed through a linear projection to produce the final output feature $\mathbf{F}_{3}$. Critically, this architecture enables single-directional scanning—unlike conventional bidirectional Mamba variants requiring multiple scans—reducing computational redundancy while preserving global coherence. This efficiency breakthrough not only accelerates inference but also facilitates seamless integration with subsequent enhancement stages, where the refined low-frequency features serve as stable anchors for high-frequency detail reconstruction.

### v2

@@ Abstract
Face super-resolution (FSR) under limited computational budgets remains challenging. Existing methods often treat all facial pixels equally, leading to suboptimal resource allocation and degraded performance. CNNs are sensitive to high-frequency facial features such as contours and outlines, while Mamba excels at capturing low-frequency attributes like facial color and texture with lower complexity than Transformers. Motivated by this, we propose FADPNet, a Frequency-Aware Dual-Path Network that decomposes facial features into low- and high-frequency components for dedicated processing. The low-frequency branch employs a Mamba-based Low-Frequency Enhancement Block (LFEB) that integrates state-space attention with squeeze-and-excitation to restore global interactions and emphasize informative channels. The high-frequency branch uses a CNN-based Deep Position-Aware Attention (DPA) module to refine structural details, followed by a lightweight High-Frequency Refinement (HFR) module for further frequency-specific refinement. These designs enable FADPNet to achieve a strong balance between FSR quality and efficiency, outperforming existing methods.
@@ I Introduction
Face super-resolution (FSR) focuses on reconstructing high-resolution (HR) facial images from low-resolution (LR) inputs, which plays a critical role in real-world scenarios where LR faces may appear in surveillance videos, historical archives, or mobile imaging. Enhancing facial image resolution is essential for improving downstream tasks such as face recognition [1] and face alignment [2].
Figure 1: (a) We divide the Fourier spectrum into low-, mid-, and high-frequency bands, and the corresponding energy ratios are averaged over the CelebA test set [3]. Results show that Mamba preserves a noticeably larger proportion of low-frequency energy, whereas CNN retains relatively stronger high-frequency responses. (b) Efficiency trade-offs on CelebA test set [3], demonstrating our method’s optimal balance between PSNR, Params, and inference speed.
@@
One core challenge in FSR is the long-tailed distribution of image frequency components [4]. High-frequency regions, such as the eyes, nose, mouth, and facial contours, occupy only a small portion of the face yet require disproportionate modeling capacity due to their variability. They are highly sensitive to changes in identity and expression, making them more difficult [5] to reconstruct than low-frequency areas like skin, which contribute little to identity-specific features. This imbalance creates a bottleneck—failure to recover high-frequency content leads to perceptual degradation. However, most existing methods treat all pixels equally, ignoring the skewed distribution of visual complexity, which results in inefficient resource use and suboptimal reconstruction. Although some works attempt frequency separation [2], they often fail to balance global and local facial structures.
To address these limitations, we aim to adaptively distinguish and process high- and low-frequency information. Meanwhile, we wonder about a dual-path feature fusion to mitigate the imbalance between local and global modeling in prior frequency-aware methods. Low-frequency facial components are mainly associated with smooth, slowly varying, and spatially correlated patterns, such as coarse facial geometry, contour continuity, illumination distribution, and skin-tone consistency [6]. These characteristics favor state-space modeling, whose recurrent state propagation enables efficient long-range context aggregation and promotes globally coherent representations [7, 8, 9]. In contrast, high-frequency facial components are concentrated around localized discontinuities and rapid spatial transitions, such as eye boundaries, mouth contours, nose edges, and fine texture details [6, 10]. Convolutions are more effective for such content because they act on local neighborhoods and respond strongly to spatial gradients and neighborhood-level variations. Consistent with this interpretation, we compare the band-wise frequency characteristics of intermediate features extracted by CNN [11] and Mamba [7]. As shown in Fig. 1(a), Mamba preserves a noticeably larger proportion of low-frequency energy, whereas CNN retains relatively stronger high-frequency responses. These observations lead us to consider an alternative design: could we develop a two-branch architecture that allows Mamba and CNN to each specialize in the frequency domains they are most suited for, while integrating their outputs to better balance global and local facial structure representation?
@@
Based on the above observations and analysis, we propose the Frequency-Aware Dual-Path Network (FADPNet), which allocates distinct submodules to different frequency domains. The Low-Frequency Enhancement Block (LFEB) integrates an Attentive State Space Block (ASSB) and Squeeze-and-Excitation Block (SEB) to restore stable, identity-preserving low-frequency features. The High-Frequency Enhancement Block (HFEB) combines a High-Frequency Refinement (HFR) module and Depthwise Position-aware Attention (DPA) to refine fine-grained, structure-critical details. These modules leverage complementary global and local modeling to capture facial characteristics across frequency levels. By decomposing and processing features according to frequency importance, our approach boosts facial fidelity and improves efficiency by focusing resources on perceptually critical regions. As shown in Fig. 1(b), FADPNet achieves superior efficiency and reconstruction quality compared to existing FSR methods. The main contributions of this paper are summarized as follows:
@@ II-A Face Super-Resolution
Face super-resolution [12] has made notable strides with the advancement of deep learning. Early approaches [13, 2, 14, 15] leveraged explicit facial priors to guide reconstruction. For instance, FSRNet [13] incorporated facial landmarks and parsing maps to enforce structural consistency, while FAN [2] applied a face alignment network with progressive training to enhance realism. DIC [14] introduced an iterative refinement loop combining facial components and landmark predictions. To improve the representation of facial spatial features, Hu et al. [15] first utilized 3D shape priors in the network to better preserve sharp facial spatial structures.
However, prior-based methods rely heavily on accurate prior estimation, which is unreliable at extremely low resolutions and adds computational overhead. To address this, recent works have explored attention-based [16] or data-driven [17] strategies. SPARNet [16] used spatial attention to focus on key facial regions, SISN [18] decoupled attention into structural and textural streams, and AD-GNN [19] combined spatial attention with graph-based message passing. LAAT [20] introduced local-aware refinement within a Transformer framework, SCTANet [21] jointly modeled spatial and channel attention, SFMNet [22] employed parallel frequency-spatial branches for multi-scale learning, and WFEN [23] used wavelet-based encoding to mitigate downsampling artifacts. Recent diffusion-based face restoration methods [24, 25], such as OSDFace [25], further demonstrated the growing interest in recovering high-quality facial structures and details with powerful generative priors. Nevertheless, most existing FSR methods apply uniform processing across spatial regions, ignoring the varying complexity of frequency components. In contrast, we propose a frequency-aware dual-path design that adaptively handles high- and low-frequency features at multiple scales, achieving more robust and detail-preserving reconstruction under LR conditions.
@@ II-B Long-Range Modeling
Modeling long-range dependencies is essential for FSR, as facial structures span distant spatial regions. Transformer-based methods [20, 22] achieve strong global modeling via self-attention. For instance, LAAT [20] refined fine-grained facial structures within a Transformer framework, SFMNet [22] combined spatial and frequency branches to capture global and local features, and CTCNet [26] fused multi-scale global representations from Transformers.
However, vision Transformers [26, 27] incur quadratic complexity, limiting scalability. This has motivated exploration of lower-cost alternatives for long-range modeling. Structured State-Space Models offer linear-time global modeling. Mamba[7] introduced a selective scan mechanism for efficient long-sequence processing with strong global context, and has been adapted to visual tasks [28]. MambaIR [8] extended Mamba to image restoration, addressing local pixel forgetting and redundant channel representations, while variants like MambaLLIE [29] show its adaptability to other low-level vision tasks. Despite its efficiency, Mamba inherits a causal bias from its sequence-oriented design, conflicting with the non-causal nature of super-resolution [8] and limiting spatial modeling. MambaIRv2 [9] alleviated this via direction-agnostic scanning for richer context modeling with linear scalability. DVMSR [30] demonstrated Vision Mamba’s linear-complexity global receptive field for efficient super-resolution, while MaIR [31] further validated its locality-preserving long-range modeling essential for low-frequency statistics. Unlike existing Mamba-based methods [8, 29, 9], we adopt a customized non-causal state-space backbone in the low-frequency branch, explicitly designed for FSR to better preserve structural consistency and global facial context.
@@ II-C Frequency-Aware Super-Resolution
Frequency-aware SR methods have proven effective for enhancing structures, particularly edges and textures [32, 33]. By decomposing image content into distinct frequency bands, they enable networks to model the characteristics of each component more precisely [34]. Low-frequency components typically represent smooth, homogeneous regions, while high-frequency components capture fine details and edges essential for perceptual quality [6].
Explicit transformation methods decompose reconstruction into frequency-specific operations. Wavelet-based methods, such as SRCliqueNet [6], used a discrete wavelet transform to split LR images into sub-bands, processed by specialized subnetworks to sharpen high-frequency components while preserving low-frequency structures. FreqNet [32] operated in the discrete cosine transform domain, predicting HR frequency coefficients with learnable spectral decoders. In contrast, implicit methods incorporate frequency awareness into network design. SRDN [10] employed high-pass constrained convolutions and residual connections to emphasize high-frequency propagation, AFD [35] estimated feature drifting in the DCT domain to adaptively enhance features. Recent works such as LSRNA [36] further improved detail fidelity in higher-resolution diffusion-based image generation through latent-space super-resolution and region-aware enhancement. However, existing frequency-aware methods [37] usually improve reconstruction through frequency enhancement or decomposition, while the different modeling needs of low- and high-frequency facial information are not explicitly addressed. In contrast, our dual-branch framework separately captures global low-frequency structures and local high-frequency details, improving both reconstruction quality and efficiency.
@@ III-A Overview of FADPNet
The proposed architecture, shown in Fig. 2, follows a hierarchical frequency-aware framework. Given an input image $\mathbf{I}_{LR}\in\mathbb{R}^{3\times H\times W}$, shallow features ${\mathbf{f}_{1}}\in\mathbb{R}^{C\times H\times W}$ are first extracted. We adopt a three-level U-Net, where each stage uses a cascaded basic block to decompose features into high- and low-frequency components for separate processing. The High-Frequency Enhancement Block (HFEB) refines fine details via convolutions, Residual Blocks (RB), a High-Frequency Refinement (HFR) module, and Depthwise Position-Aware Attention (DPA). The Low-Frequency Enhancement Block (LFEB) employs a dual-branch design: a Squeeze-and-Excitation Block (SEB) for local identity cues and an Attentive State Space Block (ASSB) for global structure modeling. Outputs are fused and refined through a Feed-Forward Network (FFN). To mitigate misalignment from down/up-sampling, an offset-based warping module aligns coarse structures with fine details, enhancing cross-scale consistency. Finally, multi-level features are fused and reduced to reconstruct the HR face image $\mathbf{I}_{SR}\in\mathbb{R}^{3\times H\times W}$.
@@ III-B Low-Frequency Enhancement Block (LFEB)
Our Low-Frequency Enhancement Block (LFEB) models smooth facial regions (forehead, cheeks, chin) with gradual intensity variations and limited identity-specific details. Given low-frequency features $\mathbf{F}_{L}\in\mathbb{R}^{C\times H\times W}$ from the frequency separation module, LFEB applies layer normalization, followed by two branches: a Squeeze-and-Excitation Block (SEB)[38] for adaptive channel calibration and an Attentive State Space Block (ASSB) for global structure modeling. A learnable fusion scheme with trainable scaling parameters balances the streams, after which a Feed-Forward Network (FFN)[7] refines the output and adds it to the block input via a residual connection. This design restores coarse facial shape and alignment while preserving photorealism.
Figure 3: Architecture of Attentive State Space Block (ASSB).
Figure 4: The architecture of (a) High-Frequency Refinement (HFR), (b) Depthwise Position-aware Attention (DPA).
@@ III-B1 Attentive State Space Block (ASSB)
where r{r} is the inner rank, T{T} is the number of prompts and d{d} is the number of hidden states in Mamba. Instance-specific prompt matrix is governed by a routing mechanism employing $\operatorname{gumbel-softmax}$ [39], which generates L{L} instance-specific prompts $\mathbf{P}_{\text{m}}\in\mathbb{R}^{L\times T}$} dynamically associated with input pixels and relevant prompts from the pool:
@@
where $\mathrm{W}_{\text{p}}$} denotes the linear layer, $\operatorname{LogSoftmax}$ refers to the logarithm of the softmax function, and gumbel-softmax is a gumbel softmax [39] operator. Then, the Semantic Guided Neighboring (SGN) unfolds this 2D feature $\mathbf{F}_{2}$ into 1D sequences for SSE. SGN-Unfold mechanism redefines the traditional spatial-to-sequential transformation by dynamically reorganizing pixels into 1D sequences based on semantic similarity, effectively mitigating the long-range decay inherent to existing methods [9]. Notably, the semantic information that the mechanism relies on during this process stems from the prompt matrix $\mathbf{P}_{\text{m}}$:
@@
$y_{i}=(\mathbf{C}+\mathbf{P})h_{i}+\mathbf{D}x_{i}$}}. (8)
Finally, an SGN-fold inversely reconstructs the semantically sorted sequence into the original 2D layout using the reverse index for precise spatial alignment. The restored feature is then linearly projected to produce the final output $\mathbf{F}_{3}$. Unlike conventional bidirectional Mamba variants, our single-directional scanning avoids multiple passes, reducing redundancy while maintaining global coherence. This improves inference speed and enables smooth integration with subsequent enhancement stages, where refined low-frequency features anchor high-frequency detail reconstruction.